OpenStudy (anonymous):

Express sin(pi/2 - x) in terms of cosx

OpenStudy (anonymous):

\[\sin ^{2}\theta + \cos ^{2}\theta = 1\]

OpenStudy (anonymous):

\[\sin \theta = \sqrt{1-\cos ^{2}x} \]

OpenStudy (anonymous):

There should be a theta beside cos^2 instead of an x. What should I do from here?

OpenStudy (anonymous):

so you know the "cofunction" identities?

OpenStudy (anonymous):

i don't think this is the right track for this one it is true that \(\sin(\frac{\pi}{2}-\theta)=\cos(\theta)\)

OpenStudy (anonymous):

you can use the addition angle formula or you can keep track of the translations of the graph from \(y=\sin(\theta)\) to \(y=\sin(\frac{\pi}{2}-\theta)\)

OpenStudy (anonymous):

easiest to use \[\sin(\alpha -\beta)=\sin(\alpha)\cos(\beta)-\sin(\beta)\cos(\alpha)\] with \[\alpha =\frac{\pi}{2},\beta=x\]

OpenStudy (anonymous):

you will get it right away

OpenStudy (anonymous):

Thats alot of trig id's, thanks! I think I can do it from here.